Tyler Tooling Company uses a job order cost system with overhead applied to products on the basis of machine hours. For the upcoming year, the company estimated its total manufacturing overhead cost at $228,660 and total machine hours at 61,800. During the first month of operations, the company worked on three jobs and recorded the following actual direct materials cost, direct labor cost, and machine hours for each job:

Job 101 Job 102 Job 103 Total
Direct materials used $ 10,800 $ 8,100 $ 5,000 $ 23,900
Direct labor $ 18,000 $ 6,200 $ 5,700 $ 29,900
Machine hours 1,400 hours 2,600 hours 800 hours 4,800 hours


Job 101 was completed and sold for $50,400.
Job 102 was completed but not sold.
Job 103 is still in process.

Actual overhead costs recorded during the first month of operations totaled $16,760.

Required:
1. Calculate the predetermined overhead rate. (Round your answer to 2 decimal places.)

2. Compute the total manufacturing overhead applied to the Work in Process Inventory account during the first month of operations. (Round your intermediate calculations to 2 decimal places.)

3. Compute the balance in the Work in Process Inventory account at the end of the first month. (Round your intermediate calculations to 2 decimal places.)

4. How much gross profit would the company report during the first month of operations before making an adjustment for over- or underapplied manufacturing overhead? (Round your intermediate calculations to 2 decimal places.)

5-a. Determine the balance in the Manufacturing Overhead account at the end of the first month. (Round your intermediate calculations to 2 decimal places.)

5-b. Is it over- or underapplied?
Answer #1

1. Predetermined overhead rate = Estimated total manufacturing overhead cost / Estimated hours = 228,660 / 61800 = 3.7 per hour

2.  Total manufacturing overhead applied to the Work in Process Inventory account = 3.7 * 4800 = 17760

3.  Balance in the Work in Process Inventory account = 5000 + 5700 + 800 * 3.7 = 13660

4. Gross profit = 50400 - 10800 - 18000 - (3.7* 1400) = 16420

5. Balance in manufacturing overhead = 17760 - 16760 = 1000

5b. It is overapplied.